La presentazione è in caricamento. Aspetta per favore

La presentazione è in caricamento. Aspetta per favore

RPC Status Outline Barrel installation and commissioning status

Presentazioni simili


Presentazione sul tema: "RPC Status Outline Barrel installation and commissioning status"— Transcript della presentazione:

1 RPC Status Outline Barrel installation and commissioning status
Trigger installation Full system commissioning - Tower tests - Global runs Gas issues Gas calibration system HV DCS-DQM Richieste

2 Full system commissioning
RPC hardware schema Full system commissioning Trigger Boards in Trigger Crates Splitter PAC GB & Sorter OPTO rec. FPGA (x 6) RMB Sorters GMT DAQ DCC S-Links Data Concentrator Card Tower rack LBBox RBC UXC area Barrel Wheel USC area

3 Installation All chamber installed except MB2 –MB3 sect 1-7 W-1, W-2
HV cabling replacement (due to wrong connectorization by firm) on going: W0, W-1 done W-2 done by mid-July W+1, W+2 cables ready to be installed by Mid-July Access for installation in August (5 weeks) All LinkBoxes installed YB+1, YB+2 Link boards and CBs, Easy crates and LV modules installed and cabled YB0 boards installation Mid-July. YB-1, YB-2, YE-1 - routing defined, all cables ready;

4 Installation Installation of Easy-crates and test of HV modules on-going in USC S1. HV multi-conductor cables (UXC-USC) installed and connectorized but W0 (Mid-July) Test done on W+1,W+2 All UX- US fibers delivered, installation YB+ YE+ middle July. Ready for feet patch panel connections (HV cables /fibers) on W+1, W+2, W0.

5 Trigger hardware Trigger System production
Tender for the TBs production: two companies, the test boards are being tested now, Components (incl. FPGAs) should be delivered by the beginning of August In September TBs should be produced and assembled Sorter Crate (full) Trigger Crate (full) DAQ & Control Crate RBC and TTU RBCs boards ready to be installed (End July) The RPC Trigger Board to be used as TTU (Technical Trigger Unit) is almost ready under the hardware point of view. A first release of VHDL code implementing cosmic muon algorithm is ready to be uploaded and tested. Development of the software tasks started and requires some working weeks (Debug in July). Aim to use for triggering purpose in August YB0 – cables routings defined, (current status ???) YB-1, YB-2, YE-1 - routing defined, 90% of cables ready; YE+-3 - routing to be defined (current status ???) 5

6 Commissioning 5 / 480 chambers
Commissioning of the final gas distributor and gas flow equalization Done Sector Test –after installation test: strip connectivity and noise rate, current stability Done W+2 W+1 W0 W-1 W-2 Chamber replacement (gas leak or HV problem) 1 2 On-site major intervention (electronics board/connectors) 5 / 480 chambers Chamber tests after cabling: Test of all LV/DCS cables Done Test after HV cable replacement ongoing on W-, finish by Mid-July W-2, W End July W+1, W+2 during access window in August-September By September all chamber on all Wheels would be tested with final cables (HV/LV/DCS cables).

7 Full system commissioning
Test and commissioning from chamber to tower : Tower test Commissioning USC electronics: Trigger and DAQ electronics Installation and test of Software: DCS, DSS , central DAQ, Trigger Commissioning from chambers up to the Link Boards (1) and commissioning of the full DAQ/Trigger chain (2) are performed in parallel in order to be ready for the full system commissioning. Strong needed of detector/software/trigger expert on site in order to fulfill the tight commissioning schedule. Common efforts of Barrel and Forward communities in order to setup global commissioning procedures.

8 Tower test Detector performance test using the final RPC/trigger system. Readout of detector data via diagnostic readouts from LBs of one tower. Strip Connectivity Test FEB – LB Connectivity Test Noise rate as a function of HV and thresholds DCS with final configuration able to handle full wheel Develop Online XDAQ application Prepare check list and documentation Bottle neck for the full wheel test: Gas flow limited (temporary gas line) Connections YBs to the USC services not done yet. Use temporary cables https://twiki.cern.ch/twiki/bin/view/CMS/RpcTowerTestsResults

9 Schedule Global run W+2, W+1, W0 complete commissioned by end-October
All towers fully commissioned by Feb08 W-1, W-2 cannot be commissioned on the surface due to lack of electronics on surface. More effort will be put on full system commissioning of YBs and YEs. W-1, W-2 tower test will start as soon as possible after lowering (aim to finish End-Nov Global run First partecipation to the GR in July 07: W+2 (10+11 corresponding to trigger sector 9) RBCs-TTU test on W+1 in August W0 (sector 10+ sector 11) Global Run in September W+1 W+2 W0 ( sectors 10 + sector 11 in each wheel) October W+1 W+2 W0 YE+ (Full wedge) November

10 Gas system status All racks (Mixer, Humidifier, Pump, Pre-distribution, Final distribution) installed and ready Analysis systems: IR analyzer (binos) for iC4H10 to be installed Relative Humidity probe installed and ready Purifier I (molecular sieve): ready, but not commissioned (about 1 week work) Purifier II(Cu/Zn filters): installed in SGX5, but not commissioned ( ~ mid July) Pending issues: Flammable detection system not yet ready. Piping heating in the shaft foreseen in the this weeks

11 Closed loop system General remarks:
There is no doubt that the RPC gas system is complex and it represents one of the key tools for the system operation At the start up some learning period will be necessary to bring the system to a stable state In 2008 low luminosity and “presumably” short running period We have one year work to setup the system and understand operation conditions Combined ATLAS CERN CMS effort Keep level of attention high, but no panic

12 Closed loop system Possible scenario:
Four months run 3 weeks out of four per months Possible scenario: Start with a 50% fresh mixture for the first month (study purification unit) Go to 30 % fresh mixtur2 for the second month Go to 10 % fresh mixture for third and fourth month Total cost at 1 vol/3h: ~ 105 kCHF

13 Gas calibration Calibration system ready next week moved to CERN.
In view of the positive opinion expressed by the INFN ATLAS-CMS review board, we have reiterated the request for a gas-chromatograph

14 Stato degli Ordini (gare 2003)
HV Stato degli Ordini (gare 2003) 2 MAO sono stati già consegnati (W+2 wheel test) Schede HV sono ferme in attesta di risolvere il problema della deriva termica modello ordinati consegnati mancanti SY1525 2 A MACISTE 3 1 A3512N - HV 80 47 33 BC A1676 7 A CRATE 34 A3000FB - FAN 14 A3486S - MAO 5 A LV 60 57

15 HV Il problema della deriva termica
Nel 2006 abbiamo riscontrato tutta una serie di problemi relativi alla lettura della corrente della schede di HV (A3513N) Da qualche giorno la CAEN è riuscita a riprodurre il problema nei suoi laboratori, lasciando le schede accese per una settimana, come da noi suggerito. Ora stanno lavorando alla risoluzione del problema che è sia di tipo software (sincronizzazione) che di tipo hardware (grounding) Entro fine luglio avremo una schema completo dei problemi e delle possibili soluzioni

16 DCS-DQM Il sistema dei DCS degli RPC è stato integrato in quello di CMS a fine giugno ed è attualmente in test a P5 I 3 servers finanziati dall’INFN stanno lavorando senza problemi a P5 G. Polese è il nuovo responsabile del sistema I databases online sono in fase di test per quanto riguarda la parte relativa al rivelatore Tutti gli istogrammi già sviluppati per il MTCC sono stati integrati nella nuova struttura ad albero della GUI (IGUANA) E’ stato poi sviluppato un sistema di warning and alarm Definizione di un set limitato di plot che possa dare una overview dell’intero sistema

17 Richieste ME

18 Richieste MOFB Spare HV/LV Sistema base Spare Spare 2007 Costo 2007
SY1527 2 1 10.036 A3009 60 6 9.000 4 18.000 A3512N 80 5 23.230 MAO 5.271 A3485 3 4.278 A3000 34 3.000 A1676 7 2.447 TOTALE

19 Richieste MOFB


Scaricare ppt "RPC Status Outline Barrel installation and commissioning status"

Presentazioni simili


Annunci Google